# Anomaly of Subsystem Symmetries in Exotic and Foliated $BF$ Theories

### v1

@@ Abstract
We study the mixed ’t Hooft anomaly of the subsystem symmetries in the exotic $BF$ theory and the foliated $BF$ theory in 2+1 dimensions, both of which are fractonic quantum field theories describing the equivalent physics. In the anomaly inflow mechanism, the ’t Hooft anomaly of the subsystem symmetries can be cancelled by combining a subsystem symmetry-protected topological (SSPT) phase in one dimension higher. In this work, we construct the exotic and foliated $BF$ theories with background gauge fields and the exotic and foliated forms of the SSPT phases using the foliated-exotic duality. In the foliated form, we see that the non-topological operator can be viewed as a symmetry-like operator. We also show that the SSPT phases with different foliation structures cancel the same anomaly. This may provide a clue to the characterization of the ’t Hooft anomaly of subsystem symmetries.
@@ 1 Introduction
The excitations with characteristic mobility originate from subsystem global symmetry, which is one of the generalized global symmetries [7], whose symmetry operator is supported on a submanifold of a certain shape in spacetime. The submanifolds are partially topological; that is, they can only be deformed in certain directions, and the symmetry operators on different submanifolds that cannot be deformed to each other are independent [8]. Fracton phases have been studied not only as lattice models (earlier studies are in [9, 5, 10, 11]), but also as effective continuum quantum field theories (QFTs) based on tensor gauge fields [12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29]. Such fractonic QFTs have discrete rotational symmetry and the tensor gauge fields respect the rotational symmetry. Related to the lattice-like spatial symmetry, fractonic QFTs exhibit UV/IR mixing; the low-energy IR physics depends on some microscopic quantities in short-distance UV physics. As a result of this phenomenon, the tensor gauge fields can have some singularities and discontinuities in certain directions [15, 16, 17]. Subsystem symmetry in fractonic QFTs is studied in the context of generalized global symmetries, and also related to multipole symmetry [30, 31], non-invertible symmetry [32] and Symmetry TFT [33].
Although tensor gauge fields can describe some fracton phases, the QFTs constructed from foliation structure also exhibit fractonic features, and such QFTs are called foliated QFTs [34, 35, 36, 37, 38, 39, 33, 30, 31, 40, 41]. [^1] Some lattice models can also be constructed by using foliation structure. They describe the foliated fracton phases [42, 43, 44, 45, 46]. A foliation is a decomposition of a manifold into an infinite number of submanifolds: leaves. In the foliated QFT context, we consider foliations where the intersections of leaves form a lattice-like structure. Foliated QFTs contain foliated gauge fields, which are considered as lower form gauge fields on the leaves, and bulk gauge fields, which interact with foliated gauge fields. In contrast to foliated QFTs, we call fractonic QFTs with tensor gauge fields exotic QFTs. Some foliated QFTs are equivalent to the corresponding exotic QFTs [36, 38, 39, 33, 30]: this is called the foliated-exotic duality. For example, the low-energy theory of the X-cube model [11] is described by both of the exotic and foliated $BF$ theory in 3+1 dimensions [17, 35], and the tensor gauge fields in the exotic $BF$ theory explicitly correspond to the foliated and bulk gauge fields in the foliated $BF$ theory [38]. [^2] These fractonic $BF$ theories are similar to the ordinary relativistic $BF$ theory [47, 48, 49] and have several analogies with it. This structure is specific to fracton phases, and is considered as a new type of duality.
@@
If the function $\alpha(A,g)$ cannot be canceled by a local counterterm of $A$ and $g$, the QFT is said to have an ’t Hooft anomaly [50]. Due to the non-invariance of the partition function, we cannot sum over the background gauge field $A$; the global symmetry cannot be gauged. In the anomaly inflow mechanism [51], an ’t Hooft anomaly in a $d$-dimensional QFT is captured by a classical field theory in $d+1$ dimensions. In relativistic QFTs, the classical field theories are called invertible field theories [52] or symmetry-protected topological (SPT) phases [53, 54]. In this case, the classification of anomalies is interpreted as the classification of SPT phases in terms of generalized cohomology [55, 56].
@@
Then, what about the case of subsystem symmetry? In some exotic theories [25] and simple foliated theories [36], anomalies of the subsystem symmetries are captured by classical field theories called subsystem symmetry-protected topological (SSPT) phases [57, 58]. However, the relation between ’t Hooft anomalies of subsystem symmetry and SSPT phases is obscure. For example, the foliation structure of the bulk SSPT phase is not canonically determined [25]. This fact obstructs the classification of anomalies of subsystem symmetry.
@@
In this paper, we discuss the $\mathbb{Z}_{N}\times\mathbb{Z}_{N}$ mixed ’t Hooft anomaly [^3] If two global symmetries can be gauged respectively, but cannot be gauged simultaneously, the system is said to have a mixed ’t Hooft anomaly. of subsystem symmetries in the exotic and foliated $BF$ theory in 2+1 dimensions [15, 38]. Although they are equivalent, it is easier to couple the exotic $BF$ theory to background tensor gauge fields and construct the 3+1d exotic SSPT phase with two simultaneous foliations that cancels the ’t Hooft anomaly of it. After constructing them, we will assume field correspondences between the exotic and foliated $BF$ theories extending the previous result without background gauge fields [38], and construct the 2+1d foliated $BF$ theory coupled to background foliated and bulk gauge fields. In the foliated $BF$ theory, we find that the non-topological operator is considered as a symmetry-like operator. Next, using the field correspondences, we will construct the foliated form of the 3+1d SSPT phase with two foliations that cancels the ’t Hooft anomaly of the foliated $BF$ theory. Finally, we will construct a bulk SSPT phase with three simultaneous foliations. In the foliated form, it is simple to construct the SSPT phase with three foliations from the SSPT phase with two foliations we have constructed. In the exotic form, on the other hand, the relation between the two SSPT phases is not manifest. Here we will construct the exotic form of the SSPT phase with three foliations via the foliated-exotic duality. This can be seen as a systematic way to construct exotic QFTs with different foliation structures. In addition, we will see that these two foliated SSPT phases are connected via a smooth deformation. This fact is considered as a hint for characterizing ’t Hooft anomalies of subsystem symmetry.
@@ 2.1 Exotic $BF$ Theory and Symmetries
where $\hat{m}^{k}$ is an $x^{k}$-dependent gauge parameter valued in integers, and $\alpha$ is a gauge parameter in the representation $\bm{1}_{0}$. The gauge parameter $\alpha$ has its own gauge transformation: $\alpha\rightarrow\alpha+2\pi n^{1}+2\pi n^{2}$, where $n^{k}$ is an $x^{k}$-dependent gauge parameter valued in integers. Due to constant parts of $\hat{m}^{k}$ and $n^{k}$, $\hat{\phi}^{12}$ and $\alpha$ can be regarded as $U(1)$-valued: $\hat{\phi}^{12}\sim\hat{\phi}^{12}+2\pi$, $\alpha\sim\alpha+2\pi$. These tensor gauge fields and parameters can have particular types of singularities and discontinuities [15, 38].
@@
so then $\Lambda_{12}$ can be written as $\Lambda_{12}=\frac{2\pi}{N}\left(\frac{1}{l^{2}}\partial_{1}n_{1}(x^{1})+\frac{1}{l^{1}}\partial_{2}n_{2}(x^{2})\right)$, where $n_{k}(x^{k})$ is an $x^{k}$-dependent single-valued function valued in integers.
@@
where $\hat{\Lambda}^{12}$ is valued in $2\pi\mathbb{Z}/N$, so $\hat{\Lambda}^{12}$ can be written as $\hat{\Lambda}^{12}=\frac{2\pi}{N}\left(n_{1}(x^{1})+n_{2}(x^{2})\right)$, where $n_{k}(x^{k})$ is an $x^{k}$-dependent single-valued function valued in integers.
@@
so then $\Lambda_{0}$ can be written as $\Lambda_{0}=\frac{2\pi}{N}\frac{1}{l^{0}}(n_{1}(x^{1})+n_{2}(x^{2}))$, where $n_{k}(x^{k})$ is an $x^{k}$-dependent single-valued function valued in integers[27].
@@ 2.3 Exotic SSPT Phase in 3+1 Dimensions
where $\hat{s}^{k}$ is an $x^{k}$-dependent gauge parameter valued in integers, and $s$ is a gauge parameter in $\bm{1}_{0}$.
@@ 3.1 Foliated $BF$ Theory and Field Correspondences
The foliated $BF$ theory in 2+1 dimensions contains two types of foliated gauge fields, which are regarded as gauge fields on the leaves, and bulk gauge fields interacting with the foliated gauge fields on each leaf [34, 35, 36, 38]. The foliated gauge fields are $U(1)$ foliated $A$-type (1+1)-form gauge fields $A^{k}\wedge dx^{k}$ and $U(1)$ foliated $B$-type zero-form gauge fields $B^{k}$. [^9] We use the words A-type and B-type used in [38]. The bulk gauge fields are $U(1)$ one-form gauge fields $a$ and $b$. The gauge transformations of the foliated gauge fields are
@@
where $\zeta^{k}\wedge dx^{k}$ is a (0+1)-form gauge parameter, $\hat{t}^{k}$ is an $x^{k}$-dependent gauge parameter valued in integers, and $\mu$ is a zero-form bulk gauge parameter. The gauge parameter $\zeta^{k}\wedge dx^{k}$ has its own gauge transformation $\zeta^{k}\wedge dx^{k}\rightarrow\zeta^{k}\wedge dx^{k}+2\pi d\xi^{k}$, where $\xi^{k}$ is an $x^{k}$-dependent gauge parameter valued in integers. The gauge transformations of the bulk gauge fields are
@@ 3.2 Coupling to the Background Foliated Gauge Fields
where $w$ is a gauge parameter valued in an integer, which is canceled by the gauge transformation of $\hat{\chi}^{k}$ mentioned later. The background gauge transformations of $C^{k}\wedge dx^{k}$, $\hat{C}^{k}$ and $\hat{c}_{12}$ are
@@
where $\hat{\chi}^{k}\,(k=1,2)$ are $x^{k}$-dependent dynamical fields valued in integers. Their dynamical gauge transformations are
@@
$\begin{split}\delta_{g}\mathcal{L}_{\text{f}}=&\frac{iN}{2\pi}\sum_{k=1}^{2}\left[(d\hat{B}^{k}-\hat{C}^{k}+b)\wedge\lambda^{k}\wedge dx^{k}+\hat{B}^{k}d\lambda^{k}\wedge dx^{k}+\hat{\lambda}^{k}(C^{k}+d\lambda^{k})\wedge dx^{k}\right.\\ &\left.-\hat{C}^{k}\wedge d(\lambda_{k}dx^{k})-(d\hat{\lambda}^{k}-\hat{\nu})\wedge d(a_{k}dx^{k}+\lambda_{k}dx^{k})\right]\\ &+\frac{iN}{2\pi}a_{0}(d\hat{\nu})_{12}\,d^{3}x+\frac{iN}{2\pi}\lambda_{0}(\hat{c}+d\hat{\nu})_{12}\,d^{3}x\\ &+\frac{iN}{2\pi}b\wedge\left(-\sum^{2}_{k=1}\lambda^{k}\wedge dx^{k}\right)-\frac{iN}{2\pi}\hat{\nu}\wedge\left(da-c-\sum^{2}_{k=1}\lambda^{k}\wedge dx^{k}\right)\\ =&\frac{iN}{2\pi}\sum_{k=1}^{2}\left[\hat{\lambda}^{k}\,C^{k}\wedge dx^{k}-(\hat{C}^{k}+d\hat{\lambda}^{k})\wedge\left\{\lambda^{k}\wedge dx^{k}+d(\lambda_{k}dx^{k})\right\}\right]+\frac{iN}{2\pi}\lambda_{0}\hat{c}_{12}\,d^{3}x\\ &+\hat{\nu}\wedge\left(c+d\lambda+\sum^{2}_{k=1}\lambda^{k}\wedge dx^{k}\right)\,.\end{split}$ (3.70)
@@ 3.3 Foliated SSPT Phase in 3+1 Dimensions
where $\hat{u}^{k}$ is an $x^{k}$-dependent gauge parameter valued in integers, $u^{k}dx^{k}$ is a (0+1)-form gauge parameter, $s$ is a zero-form gauge parameter, and $\hat{v}$ is the gauge parameter of $p$. They correspond to the dynamical gauge parameters of $\hat{\beta}^{12}$ and $(\beta_{0},\beta_{12},\beta_{3})$ as

### v2

@@ Abstract
We study the mixed ’t Hooft anomaly of the subsystem symmetries in the exotic $BF$ theory and the foliated $BF$ theory in 2+1 dimensions, both of which are fractonic quantum field theories describing the equivalent physics. In the anomaly inflow mechanism, the ’t Hooft anomaly of the subsystem symmetries can be canceled by combining a subsystem symmetry-protected topological (SSPT) phase in one dimension higher. In this work, we construct the exotic and foliated $BF$ theories with background gauge fields, and the exotic and foliated forms of the SSPT phases using the foliated-exotic duality. In the foliated form, we see that the non-topological defect that describes a fracton can be viewed as a symmetry-like operator. We also newly construct the foliated and exotic SSPT phases with different foliation structures via the foliated-exotic duality. We can show that the SSPT phases with different foliation structures cancel the same anomaly. This may provide a clue to the characterization of the ’t Hooft anomaly of subsystem symmetries.
@@ 1 Introduction
The excitations with characteristic mobility originate from subsystem global symmetry, which is one of the generalized global symmetries [7], whose symmetry operator is supported on a submanifold of a certain shape in spacetime. The submanifolds are partially topological; that is, they can only be deformed in certain directions, and the symmetry operators on different submanifolds that cannot be deformed to each other are independent [8]. Fracton phases have been studied not only as lattice models (earlier studies are in [9, 5, 10, 11]), but also as effective continuum quantum field theories (QFTs) based on tensor gauge fields [12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30]. Such fractonic QFTs have discrete rotational symmetry and the tensor gauge fields respect the rotational symmetry. Related to the lattice-like spatial symmetry, fractonic QFTs exhibit UV/IR mixing; the low-energy IR physics depends on some microscopic quantities in short-distance UV physics. As a result of this phenomenon, the tensor gauge fields can have some singularities and discontinuities in certain directions [15, 16, 17]. Subsystem symmetry in fractonic QFTs is studied in the context of generalized global symmetries, and also related to multipole symmetry [31, 32], non-invertible symmetry [33] and Symmetry TFT [34].
Although tensor gauge fields can describe some fracton phases, the QFTs constructed from foliation structure also exhibit fractonic features, and such QFTs are called foliated QFTs [35, 36, 37, 38, 39, 40, 34, 31, 32, 41, 42]. [^1] Some lattice models can also be constructed by using foliation structure. They describe the foliated fracton phases [43, 44, 45, 46, 47]. A foliation is a decomposition of a manifold into an infinite number of submanifolds: leaves. In the foliated QFT context, we consider foliations where the intersections of leaves form a lattice-like structure. Foliated QFTs contain foliated gauge fields, which are considered as lower form gauge fields on the leaves, and bulk gauge fields, which mediate the foliated gauge fields. In contrast to foliated QFTs, we call fractonic QFTs with tensor gauge fields exotic QFTs. Some foliated QFTs are equivalent to the corresponding exotic QFTs [37, 39, 40, 34, 31]: this is called the foliated-exotic duality. For example, the low-energy theory of the X-cube model [11] is described by both of the exotic and foliated $BF$ theory in 3+1 dimensions [17, 36], and the tensor gauge fields in the exotic $BF$ theory explicitly correspond to the foliated and bulk gauge fields in the foliated $BF$ theory [39]. [^2] These fractonic $BF$ theories are similar to the ordinary relativistic $BF$ theory [48, 49, 50] and have several analogies with it. This structure is specific to fracton phases, and is considered as a new type of duality.
@@
If the function $\alpha(A,g)$ cannot be canceled by a local counterterm of $A$ and $g$, the QFT is said to have an ’t Hooft anomaly [51]. Due to the non-invariance of the partition function, we cannot sum over the background gauge field $A$; the global symmetry cannot be gauged. In the anomaly inflow mechanism [52], an ’t Hooft anomaly in a $d$-dimensional QFT is captured by a classical field theory in $d+1$ dimensions. In relativistic QFTs, the classical field theories are called invertible field theories [53] or symmetry-protected topological (SPT) phases [54, 55]. In this case, the classification of anomalies is interpreted as the cobordism group classification of SPT phases[56, 57, 58, 59].
@@
Then, what about the case of subsystem symmetry? In some exotic theories [25, 30] and simple foliated theories [37], anomalies of the subsystem symmetries are captured by classical field theories called subsystem symmetry-protected topological (SSPT) phases [60, 61]. However, the relation between ’t Hooft anomalies of subsystem symmetry and SSPT phases is obscure. For example, the foliation structure of the bulk SSPT phase is not canonically determined [25, 30]. This fact obstructs the classification of anomalies of subsystem symmetry.
@@
In this paper, we discuss the $\mathbb{Z}_{N}\times\mathbb{Z}_{N}$ mixed ’t Hooft anomaly [^3] If two global symmetries can be gauged respectively, but cannot be gauged simultaneously, the system is said to have a mixed ’t Hooft anomaly. of subsystem symmetries in the exotic and foliated $BF$ theory in 2+1 dimensions [15, 39]. Although they are equivalent, it is easier to couple the exotic $BF$ theory to background tensor gauge fields and construct the 3+1d exotic SSPT phase with two simultaneous foliations that cancels the ’t Hooft anomaly of it. After constructing them, we will assume field correspondences between the exotic and foliated $BF$ theories extending the previous result without background gauge fields [39], and construct the 2+1d foliated $BF$ theory coupled to background foliated and bulk gauge fields. In the foliated $BF$ theory, we find that the non-topological defect that describes a fracton is considered as a symmetry-like operator. Next, using the field correspondences, we will construct the foliated form of the 3+1d SSPT phase with two foliations that cancels the ’t Hooft anomaly of the foliated $BF$ theory. Finally, we will construct a bulk SSPT phase with three simultaneous foliations. In the foliated form, it is simple to construct the SSPT phase with three foliations from the SSPT phase with two foliations we have constructed. In the exotic form, on the other hand, the relation between the two SSPT phases is not manifest. Here we will construct the exotic form of the SSPT phase with three foliations via the foliated-exotic duality. This can be seen as a systematic way to construct exotic QFTs with different foliation structures. In addition, we will see that these two foliated SSPT phases are connected via a smooth deformation. This fact is considered as a hint for characterizing ’t Hooft anomalies of subsystem symmetry.
@@ 2.1 Exotic $BF$ Theory and Symmetries
where $\hat{m}^{k}$ is an $x^{k}$-dependent integer-valued gauge parameter, and $\alpha$ is a gauge parameter in the representation $\bm{1}_{0}$. The gauge parameter $\alpha$ has its own gauge transformation: $\alpha\rightarrow\alpha+2\pi n^{1}+2\pi n^{2}$, where $n^{k}$ is an $x^{k}$-dependent integer-valued gauge parameter. Due to constant parts of $\hat{m}^{k}$ and $n^{k}$, $\hat{\phi}^{12}$ and $\alpha$ can be regarded as $U(1)$-valued: $\hat{\phi}^{12}\sim\hat{\phi}^{12}+2\pi$, $\alpha\sim\alpha+2\pi$. These tensor gauge fields and parameters can have particular types of singularities and discontinuities [15, 39].
@@
so then $\Lambda_{12}$ can be written as $\Lambda_{12}=\frac{2\pi}{N}\left(\frac{1}{l^{2}}\partial_{1}n_{1}(x^{1})+\frac{1}{l^{1}}\partial_{2}n_{2}(x^{2})\right)$, where $n_{k}(x^{k})$ is an $x^{k}$-dependent integer-valued and single-valued function.
@@
where $\hat{\Lambda}^{12}$ is valued in $2\pi\mathbb{Z}/N$, so $\hat{\Lambda}^{12}$ can be written as $\hat{\Lambda}^{12}=\frac{2\pi}{N}\left(n_{1}(x^{1})+n_{2}(x^{2})\right)$, where $n_{k}(x^{k})$ is an $x^{k}$-dependent integer-valued and single-valued function.
@@
so then $\Lambda_{0}$ can be written as $\Lambda_{0}=\frac{2\pi}{N}\frac{1}{l^{0}}(n_{1}(x^{1})+n_{2}(x^{2}))$, where $n_{k}(x^{k})$ is an $x^{k}$-dependent integer-valued and single-valued function[27].
@@ 2.3 Exotic SSPT Phase in 3+1 Dimensions
where $\hat{s}^{k}$ is an $x^{k}$-dependent integer-valued gauge parameter, and $s$ is a gauge parameter in $\bm{1}_{0}$.
@@ 3.1 Foliated $BF$ Theory and Field Correspondences
The foliated $BF$ theory in 2+1 dimensions contains two types of foliated gauge fields, which are regarded as gauge fields on the leaves, and bulk gauge fields mediating the foliated gauge fields on each leaf [35, 36, 37, 39]. [^9] The word “bulk” in the bulk gauge field has nothing to do with the word “bulk” in the bulk SSPT phase. The foliated gauge fields are $U(1)$ foliated $A$-type (1+1)-form gauge fields $A^{k}\wedge dx^{k}$ and $U(1)$ foliated $B$-type zero-form gauge fields $B^{k}$. [^10] We use the words A-type and B-type used in [39]. The bulk gauge fields are $U(1)$ one-form gauge fields $a$ and $b$. The gauge transformations of the foliated gauge fields are
@@
where $\zeta^{k}\wedge dx^{k}$ is a (0+1)-form gauge parameter, $\hat{t}^{k}$ is an $x^{k}$-dependent integer-valued gauge parameter, and $\mu$ is a zero-form bulk gauge parameter. The gauge parameter $\zeta^{k}\wedge dx^{k}$ has its own gauge transformation $\zeta^{k}\wedge dx^{k}\rightarrow\zeta^{k}\wedge dx^{k}+2\pi d\xi^{k}$, where $\xi^{k}$ is an $x^{k}$-dependent integer-valued gauge parameter. The gauge transformations of the bulk gauge fields are
@@ 3.2 Coupling to the Background Foliated Gauge Fields
where $w$ is an integer-valued gauge parameter, which is canceled by the gauge transformation of $\hat{\chi}^{k}$ mentioned later. The background gauge transformations of $C^{k}\wedge dx^{k}$, $\hat{C}^{k}$ and $\hat{c}_{12}$ are
@@
where $\hat{\chi}^{k}\,(k=1,2)$ are $x^{k}$-dependent integer-valued dynamical fields. Their dynamical gauge transformations are
@@
$\begin{split}\delta_{g}\mathcal{L}_{\text{f}}=&\frac{iN}{2\pi}\sum_{k=1}^{2}\left[(d\hat{B}^{k}-\hat{C}^{k}+b)\wedge\lambda^{k}\wedge dx^{k}+\hat{B}^{k}d\lambda^{k}\wedge dx^{k}+\hat{\lambda}^{k}(C^{k}+d\lambda^{k})\wedge dx^{k}\right.\\ &\left.-\hat{C}^{k}\wedge d(\lambda_{k}dx^{k})-(d\hat{\lambda}^{k}-\hat{\nu})\wedge d(a_{k}dx^{k}+\lambda_{k}dx^{k})\right]\\ &+\frac{iN}{2\pi}a_{0}(d\hat{\nu})_{12}\,d^{3}x+\frac{iN}{2\pi}\lambda_{0}(\hat{c}+d\hat{\nu})_{12}\,d^{3}x\\ &+\frac{iN}{2\pi}b\wedge\left(-\sum^{2}_{k=1}\lambda^{k}\wedge dx^{k}\right)-\frac{iN}{2\pi}\hat{\nu}\wedge\left(da-c-\sum^{2}_{k=1}\lambda^{k}\wedge dx^{k}\right)\\ =&\frac{iN}{2\pi}\sum_{k=1}^{2}\left[\hat{\lambda}^{k}\,C^{k}\wedge dx^{k}-(\hat{C}^{k}+d\hat{\lambda}^{k})\wedge\left\{\lambda^{k}\wedge dx^{k}+d(\lambda_{k}dx^{k})\right\}\right]+\frac{iN}{2\pi}\lambda_{0}\hat{c}_{12}\,d^{3}x\\ &+\frac{iN}{2\pi}\hat{\nu}\wedge\left(c+d\lambda+\sum^{2}_{k=1}\lambda^{k}\wedge dx^{k}\right)\,.\end{split}$ (3.70)
@@ 3.3 Foliated SSPT Phase in 3+1 Dimensions
where $\hat{u}^{k}$ is an $x^{k}$-dependent integer-valued gauge parameter, $u^{k}dx^{k}$ is a (0+1)-form gauge parameter, $s$ is a zero-form gauge parameter, and $\hat{v}$ is the gauge parameter of $p$. They correspond to the dynamical gauge parameters of $\hat{\beta}^{12}$ and $(\beta_{0},\beta_{12},\beta_{3})$ as